• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 04:29:04 on October23,2009

    Entry number 296213


    keyword=cavity gains restored

    About to attempt the low-current cavity test.
    We noticed the pre-gains were all zero.
    Perhaps a casualty of the power outage.
    We set them to the values we had found best on Sept 2 (halog 287555).
    See fig 1 and 2.
    (4,2,2) on 3H02 (aka Cav2)
    (4,1,1) on 3H03 (aka Cav3)
    We don't remember the post gains and sadly have no screen grabs,
    so we'll set these to 1 and see how it goes.

    FIGURE 1

    FIGURE 2